# GraspXL: Generating Grasping Motions for Diverse Objects at Scale

### v1

@@ Abstract
Human hands possess the dexterity to interact with diverse objects such as grasping specific parts of the objects and/or approaching them from desired directions. More importantly, humans can grasp objects of any shape without object-specific skills. Recent works synthesize grasping motions following single objectives such as a desired approach heading direction or a grasping area. Moreover, they usually rely on expensive 3D hand-object data during training and inference, which limits their capability to synthesize grasping motions for unseen objects at scale. In this paper, we unify the generation of hand-object grasping motions across multiple motion objectives, diverse object shapes and dexterous hand morphologies in a policy learning framework GraspXL. The objectives are composed of the graspable area, heading direction during approach, wrist rotation, and hand position. Without requiring any 3D hand-object interaction data, our policy trained with 58 objects can robustly synthesize diverse grasping motions for more than 500k unseen objects with a success rate of 82.2%. At the same time, the policy adheres to objectives, which enables the generation of diverse grasps per object. Moreover, we show that our framework can be deployed to different dexterous hands and work with reconstructed or generated objects. We quantitatively and qualitatively evaluate our method to show the efficacy of our approach. Our model and code will be available. Project page: https://eth-ait.github.io/GraspXL.
@@
Figure 2: Objective-driven Grasping Motion Synthesis. Given a hand model and an object, our goal is to synthesize grasp motions that adhere to high-level objectives, which may consist of one or multiple objectives including graspable areas (indicated by the shadow), heading directions (indicated by the red arrow), wrist rotations (indicated by the black arrow), and positions of the hand (indicated by the green dot). For each sequence, darker hand represents more recent in time.
@@ 1 Introduction
In summary, our contributions are: 1) GraspXL, a framework that synthesizes grasping motions on a large scale (500k+) of unseen objects, without relying on hand-object datasets during training. 2) A learning curriculum and objective-driven guidance to enable our method to achieve stable grasping while satisfying multiple objectives. 3) We show that our method is general enough to be deployed on reconstructed or generated objects and different dexterous hand platforms. We will release the code and models upon publication.
@@ 2 Related Work
We categorize related works into hand-object interaction synthesis and dexterous robot hand manipulation. Table 1 compares different grasping motion synthesis methods with ours regarding the use of objectives, different hand morphologies, whether they require datapoints at inference time, and the number of reported results on unseen objects.
@@ 2.1 Hand-object Interaction Synthesis
In the literature, hand interaction primarily focuses on hand(-object) reconstruction [40, 19, 26, 3, 46, 18, 14, 53, 16], static grasp synthesis [11, 42, 23, 48] and temporal hand-object motion synthesis [52, 9, 8, 2, 38, 20, 50]. This work concentrates on the latter. In synthesis, some previous methods use pure data-driven approaches to synthesize human hand manipulation sequences and rely on post-processing to enhance physical plausibility [52, 50]. Data-driven methods are supervised by 3D hand-object annotated sequences during training [52, 20, 38, 8], and some methods rely on references like wrist trajectories [50] during inference. However, acquiring accurate 3D hand-object data is infeasible to scale because of expensive capture setups [19, 39]. Due to limited training data, data-driven methods are constrained by their training distribution, making generalization challenging, especially for conditional generation tasks with motion objectives. Moreover, their dependence on data restricts their applicability across different dexterous hand platforms such as Shadow Hand [36] and Allegro Hand [44].
@@ 2.2 Dexterous Robot Hand Manipulation
Dexterous manipulation plays a crucial role in enhancing robot capabilities [49, 37, 47, 24]. With the motivation to learn from humans, some approaches use imitation learning. However, this requires full human demonstrations for both training and inference [32, 6, 25], which are usually expensive to collect. Some other methods utilize retargeted human demonstrations during training [47], use teleoperated sequences as training data [33], or learn a parameterized reward function based on demonstrations [10]. However, the reliance on expensive full grasping sequences during training still limits their generalization ability for out-of-distribution settings. Instead of relying on full trajectories, Xu et al. [45] use static pose references to guide the motion by first predicting the contact map and then optimizing the poses accordingly, with physics-based heuristics used to filter out invalid poses. This process is expensive and makes real-time inference infeasible. In contrast, our method generate motions at real-time while adhere to fine-grained motion objectives. Furthermore, without reliance on hand-object interaction data, our method can easily scale to 500k objects.
@@ 3 GraspXL
Task Definition: As illustrated in Fig. 4, we assume a hand model $h$ with $L$ links where $\mathbf{h}_{i}\in\rm I\!R^{3}$ is the position of the i-th link. The hand pose consists of joint angles $\mathbf{q}\in\rm I\!R^{L\times 3}$ and the global orientation represented by the heading direction $\mathbf{v}\in\rm I\!R^{3}$ and wrist rotation $\omega\in\rm I\!R$ about the directional vector $\mathbf{v}$. As the hand moves, it has linear and angular velocities $\mathbf{u}_{h}\in\rm I\!R^{6}$. We define $\mathbf{m\in\rm I\!R^{3}}$ as the midpoint between the thumb tip and the third joint of the middle finger. We also assume a rigid object $o$ which has a point cloud of 3D vertices $\{\mathbf{o}_{j}\}$. A user may partition the point cloud into graspable/non-graspable points. Specifically, $\{\mathbf{o}_{j}\}=\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$ where we split the points into two disjoint sets to specify whether the point should be encouraged to be in contact with the hand or not when motion is generated. As the object moves, it also has linear and angular velocities $\mathbf{u}_{o}\in\rm I\!R^{6}$. The hand links may be in contact $\mathbf{c}\in\{0,1\}^{L}$ with the object using forces in magnitude $\mathbf{f}\in\rm I\!R^{L}$. In particular, the links may contact with the graspable/non-graspable area, which are denoted as $\mathbf{c}^{+}\in\{0,1\}^{L}$ an $\mathbf{c}^{-}\in\{0,1\}^{L}$. Similarly, the force magnitude vector $\mathbf{f}$ can be decomposed into $\mathbf{f}^{+}\in\rm I\!R^{L}$ and $\mathbf{f}^{-}\in\rm I\!R^{L}$.
@@
Overview: Figure 3 outlines our RL-based method, GraspXL. Given the objectives $\mathcal{T}$ and the state s obtained from the physics simulation, we utilize a feature extraction layer $\Phi$ to derive the features. Subsequently, the policy $\boldsymbol{\pi}$ takes features as input and generates the actions a, representing the PD-control targets used to compute the torques $\boldsymbol{\tau}$. These torques are then applied to the joints of the hand model in the physics simulation to update the state s, which is subsequently fed back into our feature extraction layer for the next iteration.
@@
which includes finger joint angles q, the finger joint tracking error (compared with the target angles for PD controller) d, the hand and object velocities $\mathbf{u}_{h}$ and $\mathbf{u}_{o}$, the contact vector c, the force magnitude vector f. Finally, $\tilde{\textbf{v}},\tilde{\textbf{m}},\tilde{\omega}$ represent the differences between current and target heading directions, midpoint positions, and wrist rotation angles. To represent shape features of an object and help the hand get aware of the graspable/non-graspable areas, we construct distance features $\textbf{l}^{+}\in\rm I\!R^{L\times 3}$ where each row is the difference vector between a link position $\mathbf{h}_{i}$ to the closet object vertex in the graspable part $\{\mathbf{o}^{+}_{j}\}$. Similarly, we construct $\textbf{l}^{-}\in\rm I\!R^{L\times 3}$ for the non-graspable part $\{\mathbf{o}^{-}_{j}\}$. We ablate this component in Section 4.4.
@@
where $r_{\mathbf{c}}$ assesses the contact between fingers and the object, considering both the target and non-target areas ($r_{\mathbf{c}}=w_{c}^{+}\left\lVert\textbf{c}^{+}\right\rVert^{2}-w_{c}^{-}\left\lVert\textbf{c}^{-}\right\rVert^{2}$). The force term, $r_{\mathbf{f}}$, encourages the contact forces with graspable area and punishes the contact forces with the non-graspable area, capped by a factor proportional to the object’s weight $w_{o}$: ($r_{\mathbf{f}}=w_{f}^{+}\textbf{c}^{+}\cdot min(\textbf{f}^{+},\lambda w_{o})-w_{f}^{-}\textbf{c}^{-}\cdot min(\textbf{f}^{-},\lambda w_{o})$). $r_{\text{reg}}$ penalizes excessive velocities to ensure stability ($r_{\text{reg}}=-w_{h}||\mathbf{u}_{h}||^{2}-w_{o}||\mathbf{u}_{o}||^{2}$). If the hand model is MANO [34], we additionally apply the anatomy reward $r_{\text{anatomy}}$ [46] on joint angles $\mathbf{q}$ for natural poses.
Curriculum: Generating grasping motion with objectives requires a policy that establishes stable grasps while achieving specific goals. This is complicated because of the potential for adverse outcomes like object flipping caused by contact when trying to accomplish the objectives. To mitigate this, we introduce a learning curriculum: we start by training the policy on static objects with increased $r_{\text{goal}}$ to hone precise finger movements for objectives. Training progresses to moving objects with a higher $r_{\text{grasp}}$, enhancing wrist movements for secure, non-slip grasps. The effectiveness of this approach is discussed in Section 4.4.
Objective-driven hand guidance: Diverse objectives complicate policy exploration due to the need for varied wrist movements. To address this, we present a simple-yet-effective method to guide the hand during training and inference, improving exploration and control precision. Essentially, we compute the differences between the target and the current values for the heading direction $\mathbf{v}$, the wrist rotation angle $\omega$, and the midpoint position $\mathbf{m}$. These differences are then applied directly as bias terms for the wrist’s 6 degrees of freedom (DoF) PD-controller to guide the wrist toward motion objectives. This simple trick promotes quicker convergence and boosts performance, as detailed in Section 4.4.
@@ 4 Experiments
This section involves several experiments to evaluate our method’s effectiveness and generalization capabilities. We detail the experimental setups in Section 4.1, compare our method’s performance against others in Section 4.2, and examine its generalization across various unseen objects and hand models in Section 4.3. Lastly, we ablate our method’s components and analyze the influence of different objective combinations in Section 4.4.
@@ 4.1 Experimental Setup
Datasets: To construct the training set, we randomly select 26 objects and 32 objects from ShapeNet [4] and PartNet [29] respectively. To demonstrate our method’s generalization to large-scale object datasets, we use Objaverse [12]. Since not all objects in the three datasets are suitable for rigid-body grasping (e.g. a piece of paper), we filter out objects with such shapes. After filtering and preprocessing, we have 48, 3993 and 503k objects in PartNet, ShapeNet, and Objaverse for test, respectively. The PartNet test set contains unseen objects from categories seen during training, while the objects of the other two test sets are completely novel. Please see SupMat for details on dataset filtering and preprocessing.
@@
Evaluation Protocol: We measure the stability of the grasps with a success metric metric. To assess the adherence to the objectives, we compute several metrics and utilize the ShapeNet and PartNet datasets, which contains part-based objects. Midpoint Error (Mid. Error): The mean Euclidean distance between the final midpoint position $\mathbf{m}$ and the target $\bar{\mathbf{m}}$ measured in centimeters. Heading Error (Head. Error): The mean geodesic distance between the final heading direction $\mathbf{v}$ and the target $\bar{\mathbf{v}}$ measured in radian. Wrist Rotation Error (Rot. Error): The mean absolute error between the final wrist rotation $\omega$ and the target $\bar{\omega}$ measured in radian. Contact Ratio: The ratio between the number of links contacted with the graspable part $\{\mathbf{o}^{+}_{j}\}$ and the entire object $\{\mathbf{o}_{j}\}$. Grasping Success Rate (Suc. Rate): A grasp is determined as success if the object is lifted higher than 10cm and remains stable without falling until the sequence terminates.
Baselines: We choose SynH2R as main baseline, because it offers some degree of controllability in the motion generation, (in terms of heading direction $\mathbf{v}$), we adapt it with more fine-grained motion objectives. In particular, we have the following baselines. SynH2R: The original method generates grasping motion by first generating static grasping reference poses with an optimization procedure. Then it uses an RL-based policy to approach and follow the grasping reference pose. To adapt this method for objective-driven synthesis beyond heading direction $\mathbf{v}$, we add additional optimization terms to include the wrist rotation error $\omega$ and the midpoint error $\mathbf{m}$. We also encourage the contacts with the graspable part while punishing the contacts with the non-graspable part during the optimization. The RL-based policy stays the same as the original one in SynH2R, which is trained with the generated grasping references. SynH2R-PD: The grasping references are generated the same way as in SynH2R above. We then set the reference poses as targets for the PD controller following [9].

### v2

@@ Abstract
Human hands possess the dexterity to interact with diverse objects such as grasping specific parts of the objects and/or approaching them from desired directions. More importantly, humans can grasp objects of any shape without object-specific skills. Recent works synthesize grasping motions following single objectives such as a desired approach heading direction or a grasping area. Moreover, they usually rely on expensive 3D hand-object data during training and inference, which limits their capability to synthesize grasping motions for unseen objects at scale. In this paper, we unify the generation of hand-object grasping motions across multiple motion objectives, diverse object shapes and dexterous hand morphologies in a policy learning framework GraspXL. The objectives are composed of the graspable area, heading direction, wrist rotation, and hand position. Without requiring any 3D hand-object interaction data, our policy trained with 58 objects can robustly synthesize diverse grasping motions for more than 500k unseen objects with a success rate of 82.2%. At the same time, the policy adheres to objectives, which enables the generation of diverse grasps per object. Moreover, we show that our framework can be deployed to different dexterous hands and work with reconstructed or generated objects. We quantitatively and qualitatively evaluate our method to show the efficacy of our approach. Our model, code, and the large-scale generated motions are available at https://eth-ait.github.io/graspxl/.
@@
Figure 2: Objective-driven Grasping Motion Synthesis. Given a hand model and an object, our goal is to synthesize grasp motions that adhere to high-level objectives, which may consist of one or multiple objectives including graspable areas (indicated by the shadow), heading directions (indicated by the red arrow), wrist rotations (indicated by the black arrow), and positions of the hand (indicated by the green dot). For each sequence, the darker hand represents more recent in time.
**footnotetext: Now at HKUST(GZ)&HKUST
@@ 1 Introduction
In summary, our contributions are: 1) GraspXL, a framework that synthesizes grasping motions on a large scale (500k+) of unseen objects, without relying on hand-object datasets during training. 2) A learning curriculum and objective-driven guidance to enable our method to achieve stable grasping while satisfying multiple objectives. 3) A dataset of diverse generated grasp motions for 500k+ objects with different hands. 4) We show that our method is general enough to be deployed on reconstructed or generated objects and different dexterous hands. The code, models, and dataset are released on our project page.
@@ 2 Related Work
We categorize related works into hand-object interaction synthesis and dexterous robot hand manipulation. Tab. 1 compares different grasping motion synthesis methods with ours regarding the use of objectives, different hand morphologies, whether they require datapoints at inference time, and the number of reported results on unseen objects.
@@ 2.1 Hand-object Interaction Synthesis
In the literature, hand interaction primarily focuses on hand(-object) reconstruction [40, 19, 27, 3, 46, 18, 14, 53, 16], static grasp synthesis [11, 42, 24, 48] and temporal hand-object motion synthesis [52, 9, 2, 38, 20, 50]. This work concentrates on the latter. In synthesis, some previous methods use pure data-driven approaches to synthesize human hand manipulation sequences and rely on post-processing to enhance physical plausibility [52, 50]. Data-driven methods are supervised by 3D hand-object annotated sequences during training [52, 20, 38], and some methods rely on references like wrist trajectories [50] during inference. However, acquiring accurate 3D hand-object data is infeasible to scale because of expensive capture setups [19, 39]. Due to limited training data, data-driven methods are constrained by their training distribution, making generalization challenging, especially for conditional generation tasks with motion objectives. Moreover, their dependence on data restricts their applicability across different dexterous hand platforms such as Shadow Hand [37] and Allegro Hand [44].
@@ 2.2 Dexterous Robot Hand Manipulation
Dexterous manipulation plays a crucial role in enhancing robot capabilities [49, 47, 25]. With the motivation to learn from humans, some approaches use imitation learning. However, this requires full human demonstrations for both training and inference [33, 7, 26], which are usually expensive to collect. Some other methods utilize retargeted human demonstrations during training [47], use teleoperated sequences as training data [34], or learn a parameterized reward function based on demonstrations [10]. However, the reliance on expensive full grasping sequences during training still limits their generalization ability for out-of-distribution settings. Instead of relying on full trajectories, Xu et al. [45] use static pose references to guide the motion by first predicting the contact map and then optimizing the poses accordingly, with physics-based heuristics used to filter out invalid poses. This process is expensive and makes real-time inference infeasible. In contrast, our method generates motions in real-time while adhering to fine-grained motion objectives. Furthermore, without reliance on hand-object interaction data, our method can easily scale to 500k objects.
@@ 3 GraspXL
Task Definition: As illustrated in Fig. 4, we assume a hand model $h$ with $L$ links where $\mathbf{h}_{i}\in\rm I\!R^{3}$ is the position of the i-th link. The hand pose consists of joint angles $\mathbf{q}\in\rm I\!R^{L\times 3}$ and the global orientation represented by the heading direction $\mathbf{v}\in\rm I\!R^{3}$ and wrist rotation $\omega\in\rm I\!R$ about the directional vector $\mathbf{v}$. As the hand moves, it has linear and angular velocities $\mathbf{u}_{h}\in\rm I\!R^{6}$. We define $\mathbf{m\in\rm I\!R^{3}}$ as the midpoint between the thumb tip and the third joint of the middle finger. We also assume a rigid object $o$ which has a point cloud of 3D vertices $\{\mathbf{o}_{j}\}$. A user may partition the point cloud into graspable/non-graspable points, specifically $\{\mathbf{o}_{j}\}=\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$, where we split the points into two disjoint sets to specify whether the point should be encouraged to be in contact with the hand or not when motion is generated. As the object moves, it also has linear and angular velocities $\mathbf{u}_{o}\in\rm I\!R^{6}$. The hand links may be in contact $\mathbf{c}\in\{0,1\}^{L}$ with the object using forces in magnitude $\mathbf{f}\in\rm I\!R^{L}$. In particular, the links may contact with the graspable/non-graspable area, which are denoted as $\mathbf{c}^{+}\in\{0,1\}^{L}$ an $\mathbf{c}^{-}\in\{0,1\}^{L}$. Similarly, the force magnitude vector $\mathbf{f}$ can be decomposed into $\mathbf{f}^{+}\in\rm I\!R^{L}$ and $\mathbf{f}^{-}\in\rm I\!R^{L}$.
@@
Overview: Fig. 3 outlines our RL-based method, GraspXL. Given the objectives $\mathcal{T}$ and the state s obtained from the physics simulation, we utilize a feature extraction layer $\Phi$ to derive the features. Subsequently, the policy $\boldsymbol{\pi}$ takes features as input and generates the actions a, representing the PD-control targets used to compute the torques $\boldsymbol{\tau}$. These torques are then applied to the joints of the hand model in the physics simulation to update the state s, which is subsequently fed back into our feature extraction layer for the next iteration.
@@
which includes finger joint angles q, the finger joint tracking error (compared with the target angles for PD controller) d, the hand and object velocities $\mathbf{u}_{h}$ and $\mathbf{u}_{o}$, the contact vector c, the force magnitude vector f. Finally, $\tilde{\textbf{v}},\tilde{\textbf{m}},\tilde{\omega}$ represent the differences between current and target heading directions, midpoint positions, and wrist rotation angles. To represent shape features of an object and help the hand get aware of the graspable/non-graspable areas, we construct distance features $\textbf{l}^{+}\in\rm I\!R^{L\times 3}$ where each row is the difference vector between a link position $\mathbf{h}_{i}$ to the closet object vertex in the graspable part $\{\mathbf{o}^{+}_{j}\}$. Similarly, we construct $\textbf{l}^{-}\in\rm I\!R^{L\times 3}$ for the non-graspable part $\{\mathbf{o}^{-}_{j}\}$. We ablate this component in Sec. 4.4.
@@
where $r_{\mathbf{c}}$ assesses the contact between fingers and the object, considering both the target and non-target areas ($r_{\mathbf{c}}=w_{c}^{+}\left\lVert\textbf{c}^{+}\right\rVert^{2}-w_{c}^{-}\left\lVert\textbf{c}^{-}\right\rVert^{2}$). The force term, $r_{\mathbf{f}}$, encourages the contact forces with the graspable area and punishes the contact forces with the non-graspable area, capped by a factor proportional to the object’s weight $w_{o}$: ($r_{\mathbf{f}}=w_{f}^{+}\textbf{c}^{+}\cdot min(\textbf{f}^{+},\lambda w_{o})-w_{f}^{-}\textbf{c}^{-}\cdot min(\textbf{f}^{-},\lambda w_{o})$). $r_{\text{reg}}$ penalizes excessive velocities to ensure stability ($r_{\text{reg}}=-w_{h}||\mathbf{u}_{h}||^{2}-w_{o}||\mathbf{u}_{o}||^{2}$). If the hand model is MANO [35], we additionally apply the anatomy reward $r_{\text{anatomy}}$ [46] on joint angles $\mathbf{q}$ for natural poses.
Curriculum: Generating grasping motion with objectives requires a policy that establishes stable grasps while achieving specific goals. This is complicated because of the potential for adverse outcomes like object flipping caused by contact when trying to accomplish the objectives. To mitigate this, we introduce a learning curriculum: we start by training the policy on static objects with increased $r_{\text{goal}}$ to hone precise finger movements for objectives. Training progresses to moving objects with a higher $r_{\text{grasp}}$, enhancing wrist movements for secure, non-slip grasps. The effectiveness of this approach is discussed in Sec. 4.4.
Objective-driven hand guidance: Diverse objectives complicate policy exploration due to the need for varied wrist movements. To address this, we present a simple-yet-effective method to guide the hand during training and inference, improving exploration and control precision. Essentially, we compute the differences between the target and the current values for the heading direction $\mathbf{v}$, the wrist rotation angle $\omega$, and the midpoint position $\mathbf{m}$. These differences are then applied directly as bias terms for the wrist’s 6 degrees of freedom (DoF) PD-controller to guide the wrist toward motion objectives. This simple trick promotes quicker convergence and boosts performance, as detailed in Sec. 4.4.
@@ 4 Experiments
This section involves several experiments to evaluate our method’s effectiveness and generalization capabilities. We detail the experimental setups in Sec. 4.1, compare our method’s performance against others in Sec. 4.2, and examine its generalization across various unseen objects and hand models in Sec. 4.3. Lastly, we ablate our method’s components and analyze the influence of different objective combinations in Sec. 4.4.
@@ 4.1 Experimental Setup
Datasets: To construct the training set, we randomly select 26 objects and 32 objects from ShapeNet [4] and PartNet [30] respectively. To demonstrate our method’s generalization to large-scale object datasets, we use Objaverse [12]. Since not all objects in the three datasets are suitable for rigid-body grasping (e.g. a piece of paper), we filter out objects with such shapes. After filtering and preprocessing, we have 48, 3993 and 503k objects in PartNet, ShapeNet, and Objaverse for testing, respectively. The PartNet test set contains unseen objects from seen categories, while the objects of the other two test sets are completely novel. Please see SupMat for details on dataset filtering and preprocessing.
@@
Evaluation Protocol: We measure the stability of the grasps with a success metric metric. To assess the adherence to the objectives, we compute several metrics and utilize the ShapeNet and PartNet (which contains part-based objects) datasets. Midpoint Error (Mid. Error): The mean Euclidean distance between the final midpoint position $\mathbf{m}$ and the target $\bar{\mathbf{m}}$ measured in centimeters. Heading Error (Head. Error): The mean geodesic distance between the final heading direction $\mathbf{v}$ and the target $\bar{\mathbf{v}}$ measured in radian. Wrist Rotation Error (Rot. Error): The mean absolute error between the final wrist rotation $\omega$ and the target $\bar{\omega}$ measured in radian. textbfContact Ratio: The ratio between the number of links contacted with the graspable part $\{\mathbf{o}^{+}_{j}\}$ and the entire object $\{\mathbf{o}_{j}\}$. Grasping Success Rate (Suc. Rate): A grasp is determined as a success if the object is lifted higher than 10cm and remains stable without falling until the sequence terminates.
Baselines: We choose SynH2R as the main baseline, because it offers some degree of controllability in the motion generation, (in terms of heading direction $\mathbf{v}$). We adapt it with more fine-grained motion objectives. In particular, we have the following baselines. SynH2R: The original method generates grasping motion by first generating static grasping reference poses with an optimization procedure. Then it uses an RL-based policy to approach and follow the grasping reference pose. To adapt this method for objective-driven synthesis beyond heading direction $\mathbf{v}$, we add additional optimization terms to include the wrist rotation error $\omega$ and the midpoint error $\mathbf{m}$. We also encourage the contacts with the graspable part while punishing the contacts with the non-graspable part during the optimization. The RL-based policy stays the same as the original one in SynH2R, which is trained with the generated grasping references. SynH2R-PD: The grasping references are generated the same way as in SynH2R above. We then set the reference poses as targets for the PD controller following [9].